OpenStudy (mitu12):

draw one valid Lewis structure for each compound.Assume the atoms are arranged as drawn

OpenStudy (mitu12):

\[CH _{5}N\]

OpenStudy (mayankdevnani):

do you try or not? @mitu12

OpenStudy (mitu12):

yes i did

OpenStudy (mayankdevnani):

show me

OpenStudy (mitu12):

C: 1*4=4 H:5*1=5 N: *5=5/14

OpenStudy (mitu12):